OpenStudy (anonymous):

What is the value of n in the equation n + (n + 1) + (n + 2) = 9?

OpenStudy (cj49):

3n+3=9 3n=9-3 3n=6 n=6/3 n=2

OpenStudy (anonymous):

wow you are so good!